# Static Word Embeddings for Sentence Semantic Representation

### v1

@@ Appendix A Analysis of mGTE Embeddings
Figure 3 shows the t-SNE visualisation of sentence embeddings produced by mGTE-base at the 1st, 4th, 8th, and 12th (= last) layers of Transformer. We encode 1,000 translation pairs sampled from an English–Japanese parallel corpus; as such, ideally they should have close representations regardless of language. To obtain sentence vectors, we apply mean pooling at each layer. [^13] For the last layer, we also tried using the [[[[CLS]]]] embedding following the default pooling configuration, and got a similar result. The figure shows that the embeddings are completely separated by language at the 1st and 4th layers (just like our SWEs before PCA in Figure 1), but are less clustered at the 8th layer and completely mixed at the the last layer. [^14] As such, applying PCA did not improve mGTE’s performance. This suggests that mGTE recognises the input language at lower layers, and then captures language-agnostic features such as sentence semantics at higher layers. We also verify this hypothesis by evaluating embeddings at each layer on a translation retrieval task and observing better performance at higher layers.
@@
Notably, this result is somewhat inconsistent with the previous findings that multilingual masked language models (MLMs) like mBERT have language-specific subspaces in their embedding even at higher layers Gonen et al. (2020); Libovický et al. (2020). One possible reason for this is that mGTE is fine-tuned on various tasks to produce sentence representation, which might make it generate language-agnostic embeddings. To verify this hypothesis, we also evaluate the embeddings of mGTE’s backbone MLM (i.e. a BERT-like model with several enhancements) before it was fine-tuned on labelled data, [^15] We use the model at https://huggingface.co/Alibaba-NLP/gte-multilingual-mlm-base. and Figure 4 shows the results. It clearly indicates that the embeddings are completely separated based on language at all layers, suggesting that mGTE benefits from fine-tuning in producing language-agnostic representations. Before fine-tuning, multilingual MLMs are pre-trained to predict masked tokens in each language at the last layer, and that would naturally encourage them to have language-specific subspaces, as discussed in Gonen et al. (2020).
@@
Models Emb Size $d$ Model Path MiniLM-L6 384 sentence-transformers/all-MiniLM-L6-v2 GTE-small 384 thenlper/gte-small GTE-base 768 Alibaba-NLP/gte-base-en-v1.5 GTE-large 1,024 Alibaba-NLP/gte-large-en-v1.5 mGTE 768 Alibaba-NLP/gte-multilingual-base mGTE-MLM 768 Alibaba-NLP/gte-multilingual-mlm-base BGE 768 BAAI/bge-base-en-v1.5 SimCSE 768 princeton-nlp/sup-simcse-bert-base-uncased Nomic 768 nomic-ai/nomic-embed-text-v1.5 LLaMa3 4,096 McGill-NLP/LLM2Vec-Meta-Llama-3-8B-Instruct-mntp-supervised fastText 300 crawl-300d-2M.vec Sent2vec 700 sent2vec_wiki_bigrams WordLlama 512 wordllama-l3-supercat Model2Vec 512 potion-base-32M Model2Vec (cross-lingual) 256 M2V_multilingual_output
@@
Appendix B Word Embedding Ensembling
Given the success of ensemble learning on many NLP tasks, we try combining $L$ SWEs trained with different algorithms. For the input text $z$, we generate its ensemble embedding $\mathrm{f}(z)$ as:
$\mathrm{f}(z)$ $=\frac{1}{\sqrt{\sum_{i=1}^{L}\lambda_{i}}}\left[\frac{\sqrt{\lambda_{i}}}{\|\mathrm{f}_{i}(z)\|}\mathrm{f}_{i}(z)\right]_{i=1}^{L}~{},$ (6)
[…]
The last three rows in Table 10 show the results when we ensemble OURS with WordLlama (WL) and Model2Vec (M2V); regarding the last row 2*OURS + WL + MV, we set $\lambda_{i}$ to $2$ for OURS and $1$ for the other models to increase the contribution of OURS, the best-performing SWE. It shows that it performs the best of all models in both Avg-s2s and Avg, indicating the effectiveness of our ensembling method. Looking at the scores on each task, our ensembling method successfully combines the strengths of each model, e.g. the ensembled model substantially improves performance over OURS on Reranking and Retrieval, where WordLlama performs the best of all SWEs.
Appendix C Model Details

### v2

@@ Appendix C Analysis of mGTE Embeddings
Figure 3 shows the t-SNE visualisation of sentence embeddings produced by mGTE-base at the 1st, 4th, 8th, and 12th (= last) layers of Transformer. We encode 1,000 translation pairs sampled from an English–Japanese parallel corpus; as such, ideally they should have close representations regardless of language. To obtain sentence vectors, we apply mean pooling at each layer. [^20] For the last layer, we also tried using the [[CLS]] embedding following the default pooling configuration, and observed a similar result. The figure shows that the embeddings are completely separated by language at the 1st and 4th layers (just like our SWEs before PCA in Figure 1), but are less clustered at the 8th layer and completely mixed at the the last layer. [^21] As such, applying PCA did not improve mGTE’s performance. This suggests that mGTE recognises the input language at lower layers, and then captures language-agnostic features such as sentence semantics at higher layers. We also verify this hypothesis by evaluating embeddings at each layer on a translation retrieval task and observing better performance at higher layers.
@@
Notably, this result is somewhat inconsistent with the previous findings that multilingual masked language models (MLMs) like mBERT have language-specific subspaces in their embedding, even at near the last layer Gonen et al. (2020); Libovický et al. (2020). One possible reason for this is that mGTE is fine-tuned on various tasks to produce sentence representation, which might make it generate language-agnostic embeddings. To verify this hypothesis, we also evaluate the embeddings of mGTE’s backbone MLM (i.e. a BERT-like model with several enhancements) before it was fine-tuned on labelled data, [^22] We use the model at https://huggingface.co/Alibaba-NLP/gte-multilingual-mlm-base. and Figure 4 shows the results. It clearly indicates that the embeddings are completely separated based on language at all layers, suggesting that mGTE benefits from fine-tuning in producing language-agnostic representations. Before fine-tuning, multilingual MLMs are pre-trained to predict masked tokens in each language at the last layer, and that would naturally encourage them to have language-specific subspaces, as discussed in Gonen et al. (2020).
Appendix D Word Embedding Ensembling
Given the success of ensemble learning on many NLP tasks, we try combining $J$ SWEs trained with different algorithms. Specifically, for the input text $z$, we generate its embedding $\mathrm{f}(z)$ as:
$\mathrm{f}(z)$ $=\frac{1}{\sqrt{\sum_{i=1}^{J}\lambda_{i}}}\left[\frac{\sqrt{\lambda_{i}}}{\|\mathrm{f}_{i}(z)\|}\mathrm{f}_{i}(z)\right]_{i=1}^{J}~,$ (6)
where $\mathrm{f}_{i}(z)\in\mathbb{R}^{d_{i}}$ denotes the embedding of $z$ represented by the average of the $i$-th SWE, and $\lambda_{i}$ is a scalar hyper-parameter that controls the contribution of $\mathrm{f}_{i}(z)$ to $\mathrm{f}(z)$; $\|\cdot\|$ denotes the vector norm; and $\left[\mathrm{x}_{i}\right]_{i=1}^{J}$ denotes the concatenation of the vectors $\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{J}$. The dot product of $\mathrm{f}(z)$ exactly corresponds to the weighted average of the cosine similarity measured by each embedding model $\mathrm{f}_{i}(z)$, with the weight given by $\lambda_{i}$. For simplicity, we set $\lambda_{1}=\dots=\lambda_{J}=1$ unless otherwise stated. Notably, unlike ensembling ST models, this method does not increase computational cost very much because we can pre-concatenate the embeddings of $w\in V$ as $\mathrm{{E}}(w)=\left[\mathrm{E}_{i}(w)\right]_{i=1}^{J}$, and calculate $\mathrm{f}(z)$ simply by averaging $\mathrm{E}(w)$ and normalising the embedding within each subspace of $\mathrm{E}_{i}(w)$. [^23] On the other hand, the cost for calculating similarity and memory usage increase.
The last three rows in Table 13 show the results when we ensemble OURS with WordLlama (WL) and Model2Vec (M2V); regarding the last row 2*OURS + WL + MV, we set $\lambda_{i}$ to $2$ for OURS and $1$ for the other models to increase the contribution of OURS. It shows that it performs the best in both Avg-s2s and Avg, indicating the effectiveness of ensembling different SWEs. Looking at the scores on each task, our ensembling method successfully combines the strengths of each model, e.g. ensembling OURS with WordLlama is very effective on Reranking and Retrieval, where WordLlama performs the best of all SWEs likely due to its nature of emphasising nominal information (as discussed in Appendix B).
Models ($d$) Class. Clust. PairClass. Rerank. Retr. STS Summ. Avg-s2s Avg # Data Sets 12 11 3 4 15 10 1 33 56 Sentence Transformers (STs) MiniLM-L6 (384) 63.06 42.35 82.37 58.04 41.95 78.90 30.81 65.95 56.26 SimCSE-supervised (768) 67.32 33.43 73.68 47.54 21.82 79.12 31.17 62.75 48.86 BGE-base (768) 75.53 45.77 86.55 58.86 53.25 82.40 31.07 71.18 63.55 GTE-base (768) 77.17 46.82 85.33 57.66 54.09 81.97 31.17 71.51 64.11 LLM2Vec-LLaMa3 (4,096) 75.92 46.45 87.80 59.68 56.63 83.58 30.94 72.94 65.01 Static Word Embeddings (SWEs) fastText (300) 59.09 30.75 65.77 43.63 22.21 62.82 30.33 53.68 43.05 Sent2vec (700) 61.15 31.75 71.88 47.31 27.79 65.80 29.51 56.33 46.29 WordLlama (512) [WL] 60.85 36.16 73.72 52.82 33.14 72.47 30.83 60.40 50.23 Model2Vec* (512) [M2V] 66.23 35.29 77.89 50.92 32.13 74.22 29.78 62.37 51.33 OURS (256) 67.29 36.67 78.99 50.91 31.32 75.87 30.55 63.76 51.97 OURS (512) 67.96 37.14 79.00 51.07 32.01 75.59 30.35 64.06 52.35 Ensemble of SWEs OURS + WL (1,024) 66.71 37.56 79.18 52.82 34.13 75.63 30.04 64.10 52.87 OURS + WL + MV (1,536) 67.70 37.06 79.37 52.60 34.52 75.63 30.20 64.17 53.09 2*OURS + WL + M2V (1,536) 68.17 37.60 79.54 52.39 34.32 75.89 30.22 64.55 53.28
Table 13: Results on MTEB data sets. The last three rows denote the scores when we ensemble OURS with WordLlama [WL] and Model2Vec [M2V] using the method described in Appendix D. The best scores within each subtable are bold-faced. *The scores for Model2Vec in Retr. and Avg are slightly lower than the reported scores, due to a small bug in their MS MARCO evaluation.
@@ Appendix D Word Embedding Ensembling
Models Emb Size $d$ Model Path MiniLM-L6 384 sentence-transformers/all-MiniLM-L6-v2 GTE-small 384 thenlper/gte-small GTE-base 768 Alibaba-NLP/gte-base-en-v1.5 GTE-large 1,024 Alibaba-NLP/gte-large-en-v1.5 mGTE 768 Alibaba-NLP/gte-multilingual-base BGE 768 BAAI/bge-base-en-v1.5 SimCSE 768 princeton-nlp/sup-simcse-bert-base-uncased Nomic 768 nomic-ai/nomic-embed-text-v1.5 LLaMa3 4,096 McGill-NLP/LLM2Vec-Meta-Llama-3-8B-Instruct-mntp-supervised fastText 300 crawl-300d-2M.vec Sent2vec 700 sent2vec_wiki_bigrams WordLlama 512 wordllama-l3-supercat Model2Vec (English) 512 potion-base-32M Model2Vec (cross-lingual) 256 M2V_multilingual_output